OpenStudy (anonymous):

How do I proceed to solve this equation? sin(x)-cos^2(x)=0

OpenStudy (freckles):

hint: cos^2(x)=1-sin^2(x)

OpenStudy (freckles):

\[\sin(x)-(1-\sin^2(x))=0 \\ \sin^2(x)+\sin(x)-1=0\] you have a quadratic in terms of sin(x) to solve

OpenStudy (anonymous):

Many thanks! This is one of the properties of cos^2 I was not aware of.

OpenStudy (freckles):

It is a Pythagorean identity. sin^2(x)+cos^2(x)=1